The Giscome Corporations purchases from suppliers in a quarter are equal to 75 percent of the next quarters forecost sales.
Answer #1

60 day payable period means current quarter payable will be (60)/90 = 2/3 of current quarter sales is to be paid in current month while remaining 1/3 of next quarter sales is to be paid in this month

Since 75% of purchases are made

Payables each period = 2/3(0.75) times current sales + 1/3(0.75) next period sales

Q1 = 2/3*0.75*1970+1/3*0.75*2315 =1564

Q2 =2/3*0.75*2315+1/3*0.75*1930 =1640

Q3=2/3*0.75*1930+1/3*0.75*1740 =1400

Q4=2/3*0.75*1740+1/3*0.75*2380 =1465

А F 1 Purchases 2 Payable period 3 Wages, taxes, other expenses 4 Long term financing expenses 5 Next year Q1 sales Tal B D E

1 Purchases 2 Payable period 3 Wages, taxes, other expenses 4 Long term financing expenses 5 Next year Q1 sales 90 6 0.75 nex